George has been selling 6,000 t-shirts per month for $8.00 when he increased the price to $10,00 he only sold 4,000 t-shirts . What is the best approximates the price elasticity of demand?
1
Expert's answer
2020-09-21T08:10:58-0400

The price elasticity of demand is:

"Ed = \frac{4,000 - 6,000} {10 - 8} ×\frac{10 + 8} {4,000 + 6,000} = 1.8."

So, the demand is elastic.